# Stabilizing the Splits through Minimax Decision Trees

### v2

@@ 5 Numerical experiments and applications
In this section, we provide experiments that verify the advantages of MinimaxSplit and cyclic MinimaxSplit methods in low-dimensional domains, as well as integrated ensemble methods applied to approximating a real-valued function and denoising images. Performance metrics are defined in Appendix D.1, and further numerical experiments (including those in high dimensions) are provided in Appendix D.
@@ 5.1 Avoiding end-cut preference for stable cuts
To show that the ECP is remedied in MinimaxSplit, we consider a special case of the location regression model in (Cattaneo et al., 2022), similar to the discussions in Example 1. In our experiment, we generate independent replicates with $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$ and $Y$ following the standard normal or Student’s $t(\cdot)$ distributions (pure noise), sort by $X$, and scan all possible splits in the experiments in Figure 5. We also provide two baselines: Uniform random cutting picks a split point entirely at random within the current predictor range, ensuring that there is no systematic bias yet, ignoring any data-driven signal. Sampling an observed value similarly randomizes split location among actual samples. Both serve as baseline benchmarks: they eliminate ECP, but do not offer an impurity-guided improvement mechanism.
@@
Figure 5: Empirical distributions of sample size $n=500$ over $1000$ replicates of the smaller-child proportion $\min(n_{L},n_{R})/n$ when splitting pure-noise data generated by standard normal (left), $t(3)$ (middle), and $t(1)$ (right) distributions. We also use dashed lines to denote the threshold 0.05.
Figure 6: Average leaf size (mean number of samples per leaf) versus maximum tree depth for four splitting criteria (VarianceSplit (CART), MinimaxSplit, C1, C2), at two noise levels. Error bars show one standard deviation across repeated fits. The vertical axis is on a log scale to expose imbalance: small and collapsing leaf means with low variability (as for MinimaxSplit and, to a lesser extent, CART) indicate balanced partitioning, whereas the slower decay and large spread for C1/C2 reflect persistent ECP that peels off tiny fragments and leaves dominant branches. The signal function is given by (33). Increasing the centered Gaussian noise from $\sigma=0.01$ to $\sigma=0.1$ slightly attenuates rates but does not erase the structural contrast.
@@
where $L$ is the number of leaves and $n_{j}$ is the size of leaf $j$. Small average leaf node sizes and collapsed error bars signal balanced partitions; large means with huge variance indicate persistent end-cuts leaving one dominant branch and tiny fragments. Our main methodological contribution, the MinimaxSplit algorithm can be used with ensembles and is applicable in real data, but also yields a theoretically stronger version of Buja and Lee (2001)’s formula C1 and C2, where they showed that the minimum variance (a.k.a., one-sided purity) approach can identify extreme subsets much better than traditional CART. Figure 6 records the leaf distributions as the maximum depth of the tree grows, with $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$ and signal
@@ 5.2 Fixed-horizon EEG amplitude regression
The results align with the theoretical expectations. As the tree depth increases, CART and the sum-of-variances baseline increasingly display end-cut behavior, producing piecewise constant fits that flatten toward the edges of the interval and degrade extrapolation between local extrema. MinimaxSplit, by contrast, maintains near-balanced children throughout the path, which yields visibly tighter fits to the oscillatory structure of the signal and systematically lower errors. In addition, MinimaxSplit reduces error and raises $R^{2}$ markedly as depth increases, while CART and the VarianceSplit baseline improve more slowly. The trend is most pronounced in the random holdout regime where interpolation dominates; here, the additional leaves available to deeper trees are useful only if the split rule does not collapse sample sizes, and the MinimaxSplit criterion preserves effective sample size by construction. This observation matches our finding that MinimaxSplit concentrates the smaller-child proportion away from zero and one across noise laws, thereby eliminating pathological small-node failures observed under impurity-based CART.
@@ 5.3 Image denoising with ensembles
Figure 8: Comparative analysis of various random forest (depth $K=10$, 50 estimators) approaches for image denoising using different splitting strategies. The experiment compares six random forest configurations on a noisy astronaut image: original noiseless image; noisy input; three greedily splitting methods ($m_{\mathrm{try}}=d$) using consistent splitting criteria across all weak learners (trees) — VarianceSplit, MinimaxSplit, and Cyclic MinimaxSplit; three random-dimension methods where feature selection is randomized at each node (i.e., $m_{\mathrm{try}}=1$) while location selection follows the three different criteria, as described in Section 3.2.
@@
Another notable experiment involves applying decision trees to denoise noisy images (Luo et al., 2024). We apply the different tree variants to predict pixel values based on their locations, effectively treating this as a regression problem. The experiments shown in Figures 8 and 9 compare decision tree and random forest regression methods for denoising the classic Astronaut (Van der Walt et al., 2014) grayscale image, which has been preprocessed to 256 by 256.
The goal is to recover a clean grayscale image $f:[0,1]^{2}\to[0,1]$ from noisy observations $Y=f(\mathbf{X})+\varepsilon$ at pixel centers $\mathbf{X}=(X_{1},X_{2})$ laid out on a $256\times 256$ grid (equipped with the uniform distribution). We normalize intensities to $[0,1]$ and add i.i.d. Gaussian noise $\varepsilon\buildrel\mathrm{law}\over{\sim}\mathcal{N}(0,\,0.1^{2})$. Each pixel becomes one regression sample, and we fit axis-aligned trees that approximate the unknown $f$ as a piecewise constant function on a dyadic partition, which is precisely the regime covered by our analysis of axis-aligned trees and their ensembles.
@@
Figure 8 shows qualitative reconstructions for the six forests. The VarianceSplit forest leaves noticeable grain in flat regions and blurs edges at depth $10$. In contrast, both Minimax ensembles suppress speckle while preserving high-contrast features such as facial contours and texture edges. Further, across all metrics, the MinimaxSplit forests outperforms the VarianceSplit forest, with the largest gains in SSIM, indicating better structural fidelity. The improvements are coherent with Theorem 9: the worst-child control given by the MinimaxSplit rule prevents high-variance children from persisting across levels. The single-tree performance is shown in Figure 9.
@@ 6 Conclusion
In this study, we reexamine the well-known phenomenon of ECP, which is associated with traditional CART decision trees Breiman et al. (1984). To remedy this undesirable behavior for single decision trees Ishwaran and Kogalur (2010), we introduce a novel decision tree splitting strategy, the MinimaxSplit algorithm, along with its multivariate variant, the cyclic MinimaxSplit algorithm. Unlike traditional VarianceSplit methods, which aim to reduce overall variance in the child nodes, our MinimaxSplit algorithm seeks to minimize the maximum variance within the split partitions, thereby reducing the risk of overly biased partitions. The cyclic MinimaxSplit algorithm further ensures that each dimension is used in a balanced manner throughout the tree construction process, avoiding dominance by a subset of features. We do not claim that this new method could outperform the classical CART-like splitting under all circumstances, yet we point out that our construction avoids ECP and attains a better convergence rate guarantee as the tree progresses deeper. The (non-cyclic) MinimaxSplit construction features the anti-symmetry-breaking preference—consistency is not guaranteed in certain symmetric cases, which can, however, be alleviated by using ensemble methods.
A key theoretical methodological development of this study is that the cyclic MinimaxSplit algorithm achieves an exponential MSE decay rate given sufficient data samples, without requiring additional variance decay assumptions (Theorem 6) in regression and classification settings. Furthermore, we derive empirical risk bounds for this method, establishing its robustness in different settings (Theorem 7). Using the same powerful techniques used for trees, we prove novel results on the convergence rates of univariate partition-based martingale approximations as a by-product (see Appendix A), which are of their own interest.
In addition to single-tree regression analysis, we explore ensemble learning approaches, where we combine decision trees constructed using MinimaxSplit, cyclic MinimaxSplit, and VarianceSplit methods within a random forest framework. The results demonstrate that a hybrid approach leveraging different splitting techniques yields superior performance, particularly when dealing with non-uniform data distributions (Section 3). These findings suggest that the MinimaxSplit approach can provide more stable and adaptive decision trees compared to traditional methods, especially in low-dimensional settings.
@@ Appendix A On partition-based martingale approximations
In this section, we develop the framework of partition-based martingale approximations and show that a number of constructions feature exponential convergence rates. These problems are of their own interest (see the end of Appendix A.1) and have implications in the regression tree setting. The main takeaway is that:
@@ A.1 Basic concepts
Consider nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}$ and a real-valued random variable $U$ with a finite second moment. We use the abbreviation $\sigma(\pi_{k})$ to denote the $\sigma$-algebra generated by events of the form $\{U\in A\}_{A\in\pi_{k}}$. Define $M_{k}:=\mathbb{E}[U\mid\sigma(\pi_{k})]$. If $\Pi:=\{\pi_{k}\}_{k\geqslant 0}$ is a nested sequence of partitions, $\{\sigma(\pi_{k})\}_{k\geqslant 0}$ is a filtration generated by indicator functions of sets in $\pi_{k}$’s (Doob, 1953). It follows from the tower property of conditional expectations that the sequence $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ is a martingale (in fact, a Doob martingale (Doob, 1940, 1953)). We call this martingale the $\Pi$-based martingale approximation of the random variable $U$, or in general a partition-based martingale approximation that approximates the random variable $U$ in $L^{2}$.
@@
In turn, when applying recursively each splitting rule starting from $\pi_{0}=\{\mathbb{R}\}$, we obtain a nested sequence of partitions $\Pi=\{\pi_{k}\}_{k\geqslant 0}$ depending on the law of $U$. We call the resulting $\Pi$-based martingale approximation $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ the variance (resp. Simons, minimax, median) martingale (with respect to $U$). See Figure 10 for examples of the four martingale approximations and their convergence rates.
@@
If $U$ is uniformly distributed on a compact interval, all four martingales coincide. For example, if $U\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$, each of the four martingales from Definition 14 satisfy $M_{k}=\mathbb{E}[U\mid\sigma(\pi_{k})]$, where $\pi_{k}:=\{[j/2^{k},(j+1)/2^{k}):0\leqslant j<2^{k}\}$ and it follows that $\mathbb{E}[(U-M_{k})^{2}]=4^{-k}/12$.
@@
The terminology martingale approximation has been used extensively in the probability literature with different meanings. In Rüschendorf (1985), it refers to the best approximation of a random vector by a (single) martingale based on ideas from optimal transport. In Borovskikh and Korolyuk (1997) and Hall and Heyde (2014), it refers to techniques from martingale theory (such as inequalities and CLT rates) with various applications in statistics. Similar techniques are also used in the study of stationary ergodic sequences (Wu and Woodroofe, 2004; Zhao and Woodroofe, 2008) and Markovian walks (Grama et al., 2018). Note that in our setting, $Y$ is a sum of martingale differences, but whose variance is bounded, and hence one cannot apply the martingale CLT or its convergence rates.
@@ A.2 Uniform convergence rates
    If $\{M_{k}\}_{k\geqslant 0}$ is the variance martingale, $\mathbb{E}[(U-M_{k})^{2}]\leqslant 2.71\cdot 2^{-2k/3}$.
@@
The general case of a bounded $U$ can be derived by a scaling argument, since the constructions in Definition 14 are scale-invariant. Finding the optimal rate $r\in(0,1)$ (where $\mathbb{E}[(U-M_{k})^{2}]\leqslant Lr^{k}$ for some universal constant $L$) is also an intriguing question. Observe that one cannot have $r<1/4$ by Example 15 above. In Examples 17 and 18 below, we show that the rate $r$ in Theorem 16 is indeed optimal for the Simons and median martingales.
@@
$=s^{-(J+1)}\sum_{\ell=J}^{\infty}s^{\ell+1}(2s-1)\bigg{(}1+\Big{(}\frac{1-s}{s}\Big{)}^{\ell}\bigg{)}$
$=\frac{s^{-(J+1)}(2s-1)s^{J+1}}{1-s}+\frac{s^{-(J+1)}(2s-1)s}{1-(1-s)}$
@@
$\mathbb{P}\Big{(}U=\sum_{j=1}^{k-1}s^{j-1}-s^{k-1}\Big{)}=2^{-k},\leavevmode\nobreak\ k\geqslant 1.$

### v3

@@ 5 Numerical experiments and applications
In this section, we provide experiments that verify the advantages of MinimaxSplit and cyclic MinimaxSplit methods in low-dimensional domains, as well as integrated ensemble methods applied to approximating a real-valued function and denoising images. Performance metrics are defined in Appendix E.1, and further numerical experiments (including those in high dimensions) are provided in Appendix E.
@@ 5.1 Avoiding end-cut preference for stable cuts
To show that the ECP is remedied in MinimaxSplit, we consider a special case of the location regression model in (Cattaneo et al., 2022), similar to the discussion in Example 2. In our experiment, we generate independent replicates with $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$ and $Y$ following the standard normal or Student’s $t$ distributions (pure noise), sort by $X$, and scan all possible splits in the experiments in Figure 5. We also provide two baselines: Uniform random cutting picks a split point entirely at random within the current predictor range, ensuring that there is no systematic bias, yet ignoring any data-driven signal. Sampling an observed value similarly randomizes split location among actual samples. Both serve as benchmarks: they eliminate ECP, but do not offer an impurity-guided improvement mechanism.
@@
Figure 5: Empirical CDFs of the smaller child proportion $\min(n_{L},n_{R})/n$ for sample size $n=500$, based on $1000$ replicates when splitting pure-noise data generated by standard normal (left), $t(3)$ (middle), and $t(1)$ (right) distributions. We also use dashed lines to denote the threshold 0.05.
Figure 6: Average leaf size (mean number of samples per leaf) versus maximum tree depth for six splitting criteria (VarianceSplit (CART), MinimaxSplit, C1, C2 (Buja and Lee, 2001), CovRT (Zhang and Ma, 2026), and alpha-CART Wager and Walther (2015)) at two noise levels. Error bars show one standard deviation across repeated fits. The vertical axis is on a log scale to expose imbalance: small and collapsing leaf means with low variability (as for MinimaxSplit and, to a lesser extent, CART) indicate balanced partitioning, whereas the slower decay and large spread for C1/C2 reflect persistent ECP that peels off tiny fragments and leaves dominant branches. The signal function is given by (34). Increasing the centered Gaussian noise from $\sigma=0.01$ to $\sigma=0.1$ slightly attenuates rates but does not erase the structural contrast.
@@
where $L$ is the number of leaves and $n_{j}$ is the size of leaf $j$. Small average leaf node sizes and collapsed error bars signal balanced partitions; large means with huge variance indicate persistent end-cuts leaving one dominant branch and tiny fragments. Our main methodological contribution, the MinimaxSplit algorithm, can be used with ensembles and is applicable to real data. It is also related in spirit to the C1 and C2 criteria of Buja and Lee (2001), which emphasize one-sided purity and can better identify extreme subsets than traditional CART. Figure 6 records the leaf distributions as the maximum depth of the tree grows, with $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$ and signal
@@ 5.2 Fixed-horizon EEG amplitude regression
The results align with the theoretical expectations. As the tree depth increases, CART and the sum-of-variances baselines increasingly display end-cut behavior, producing piecewise constant fits that flatten toward the edges of the interval and degrade extrapolation between local extrema. MinimaxSplit, by contrast, maintains near-balanced children throughout the path, which yields visibly tighter fits to the oscillatory structure of the signal and systematically lower errors. In addition, MinimaxSplit reduces error and raises $R^{2}$ markedly as depth increases, while CART and the VarianceSplit baseline improve more slowly. The trend is most pronounced in the random holdout regime where interpolation dominates; here, the additional leaves available to deeper trees are useful only if the split rule does not collapse sample sizes, and the MinimaxSplit criterion preserves effective sample size by construction. This observation matches our finding that MinimaxSplit concentrates the smaller-child proportion away from zero and one across noise laws, thereby eliminating pathological small-node failures observed under impurity-based CART.
@@ 5.3 Image denoising with ensembles
Figure 8: Comparative analysis of various random forest (depth $K=10$, 50 estimators) approaches for image denoising using different splitting strategies. The experiment compares six random forest configurations on a noisy astronaut image: original noiseless image; noisy input; three greedy splitting methods ($m_{\mathrm{try}}=d$) using consistent splitting criteria across all weak learners (trees) — VarianceSplit, MinimaxSplit, and Cyclic MinimaxSplit; three random-dimension methods where feature selection is randomized at each node (i.e., $m_{\mathrm{try}}=1$) while location selection follows the three different criteria, as described in Section 3.2.
@@
Another notable experiment involves applying decision trees to denoise images (Luo et al., 2024). We apply the different tree variants to predict pixel values based on their locations, effectively treating this as a regression problem. The experiments shown in Figures 8 and 9 compare decision tree and random forest regression methods for denoising the classic Astronaut (Van der Walt et al., 2014) grayscale image, which has been preprocessed to a $256\times 256$ image.
The goal is to recover a clean grayscale image $f:[0,1]^{2}\to[0,1]$ from noisy observations $Y=f(\mathbf{X})+\varepsilon$ at pixel centers $\mathbf{X}=(X_{1},X_{2})$ laid out on a $256\times 256$ grid (equipped with the uniform distribution). We normalize intensities to $[0,1]$ and add i.i.d. Gaussian noise $\varepsilon\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathcal{N}(0,\,0.1^{2})$. Each pixel becomes one regression sample, and we fit axis-aligned trees that approximate the unknown $f$ as a piecewise constant function on a dyadic partition, which is precisely the regime covered by our analysis of axis-aligned trees and their ensembles.
@@
Figure 8 shows qualitative reconstructions for the six forests. The VarianceSplit forest leaves noticeable grain in flat regions and blurs edges at depth $10$. In contrast, both Minimax ensembles suppress speckle while preserving high-contrast features such as facial contours and texture edges. Further, across all metrics, the MinimaxSplit forests outperform the VarianceSplit forest, with the largest gains in SSIM, indicating better structural fidelity. The improvements are coherent with Theorem 10: the worst-child control given by the MinimaxSplit rule prevents high-variance children from persisting across levels. The single-tree performance is shown in Figure 9.
@@ 6 Conclusion
In this study, we reexamine the well-known phenomenon of ECP, which is associated with traditional CART decision trees Breiman et al. (1984). To remedy this undesirable behavior for single decision trees Ishwaran and Kogalur (2010), we introduce a novel decision tree splitting strategy, the MinimaxSplit algorithm, along with its multivariate variant, the cyclic MinimaxSplit algorithm. Unlike traditional VarianceSplit methods, which aim to reduce overall variance in the child nodes, our MinimaxSplit algorithm seeks to minimize the maximum variance within the split partitions, thereby reducing the risk of overly biased partitions. The cyclic MinimaxSplit algorithm further ensures that each dimension is used in a balanced manner throughout the tree construction process, avoiding dominance by a subset of features. We do not claim that this new method could outperform the classical CART-like splitting under all circumstances, yet we point out that our construction avoids ECP and attains a better convergence rate guarantee as the tree grows deeper. The (non-cyclic) MinimaxSplit construction features the anti-symmetry-breaking preference—consistency is not guaranteed in certain symmetric cases, which can, however, be alleviated by using ensemble methods.
A key theoretical methodological development of this study is that the cyclic MinimaxSplit algorithm achieves an exponential MSE (or logistic loss) decay rate given sufficient data samples, without requiring additional variance decay assumptions in regression and classification settings (Theorems 7 and 13). Furthermore, we derive empirical risk bounds for this method, establishing its robustness in different settings (Theorems 8 and 14). Using the same powerful techniques used for trees, we prove novel results on the convergence rates of univariate partition-based martingale approximations as a by-product (see Appendix A), which are of independent interest.
In addition to single-tree regression analysis, we explore ensemble learning approaches, where we combine decision trees constructed using MinimaxSplit, cyclic MinimaxSplit, and VarianceSplit methods within a random forest framework. We also demonstrate a hybrid approach with different splitting techniques at different layers when dealing with non-uniform data distributions (Section E.2). These findings suggest that the MinimaxSplit approach can provide more stable and adaptive decision trees compared to traditional methods, especially in low-dimensional settings. In future work, we hope to introduce changes in the cut family and search scope (Ishwaran, 2026) to enhance its high-dimensional behavior.
@@ Appendix A On partition-based martingale approximations
In this section, we develop the framework of partition-based martingale approximations and show that a number of constructions feature exponential convergence rates. These problems are of independent interest (see the end of Appendix A.1) and have implications in the regression tree setting. The main takeaway is that:
@@ A.1 Basic concepts
Consider nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}$ and a real-valued random variable $U$ with a finite second moment. We use the abbreviation $\sigma(\pi_{k})$ to denote the $\sigma$-algebra generated by events of the form $\{U\in A\}_{A\in\pi_{k}}$. Define $M_{k}:=\mathbb{E}[U\mid\sigma(\pi_{k})]$. If $\Pi:=\{\pi_{k}\}_{k\geqslant 0}$ is a nested sequence of partitions, $\{\sigma(\pi_{k})\}_{k\geqslant 0}$ is a filtration generated by indicator functions of sets in the partitions $\{\pi_{k}\}_{k\geqslant 0}$ (Doob, 1953). It follows from the tower property of conditional expectations that the sequence $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ is a martingale (in fact, a Doob martingale (Doob, 1940, 1953)). We call this martingale the $\Pi$-based martingale approximation of the random variable $U$, or in general a partition-based martingale approximation that approximates the random variable $U$ in $L^{2}$.
@@
In turn, when applying each splitting rule recursively starting from $\pi_{0}=\{\mathbb{R}\}$, we obtain a nested sequence of partitions $\Pi=\{\pi_{k}\}_{k\geqslant 0}$ depending on the law of $U$. We call the resulting $\Pi$-based martingale approximation $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ the variance (resp. Simons, minimax, median) martingale (with respect to $U$). See Figure 10 for examples of the four martingale approximations and their convergence rates.
@@
If $U$ is uniformly distributed on a compact interval, all four martingales coincide. For example, if $U\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$, each of the four martingales from Definition 16 satisfies $M_{k}=\mathbb{E}[U\mid\sigma(\pi_{k})]$, where $\pi_{k}:=\{[j/2^{k},(j+1)/2^{k}):0\leqslant j<2^{k}\}$ and it follows that $\mathbb{E}[(U-M_{k})^{2}]=4^{-k}/12$.
@@
The terminology martingale approximation has been used extensively in the probability literature with different meanings. In Rüschendorf (1985), it refers to the best approximation of a random vector by a (single) martingale based on ideas from optimal transport. In Borovskikh and Korolyuk (1997) and Hall and Heyde (2014), it refers to techniques from martingale theory (such as inequalities and CLT rates) with various applications in statistics. Similar techniques are also used in the study of stationary ergodic sequences (Wu and Woodroofe, 2004; Zhao and Woodroofe, 2008) and Markovian walks (Grama et al., 2018). Note that in our setting, $U$ is a sum of martingale differences, but whose variance is bounded, and hence one cannot apply the martingale CLT or its convergence rates.
@@ A.2 Uniform convergence rates
    If $\{M_{k}\}_{k\geqslant 0}$ is the variance martingale, $\mathbb{E}[(U-M_{k})^{2}]\leqslant 2^{-2-2k/3}$.
@@
The general case of a bounded $U$ can be derived by a scaling argument, since the constructions in Definition 16 are scale-invariant. Finding the optimal rate $r\in(0,1)$ (where $\mathbb{E}[(U-M_{k})^{2}]\leqslant Lr^{k}$ for some universal constant $L$) is also an intriguing question. Observe that one cannot have $r<1/4$ by Example 17 above. In Examples 19 and 20 below, we show that the rate $r$ in Theorem 18 is indeed optimal for the Simons and median martingales.
@@
$=s^{-(J+1)}\sum_{\ell=J}^{\infty}s^{\ell+1}\frac{(1-s)^{2}}{2s-1}\bigg(1-\Big(\frac{1-s}{s}\Big)^{\ell}\bigg)$
$=\frac{1-s}{2s-1}\bigg(1-\Big(\frac{1-s}{s}\Big)^{J+1}\bigg)$
@@
$\mathbb{P}\Big(U=\sum_{j=1}^{k-1}s^{j-1}-s^{k-1}\Big)=2^{-k},\penalty 10000\ k\geqslant 1.$